UT 23 May 2019

Rollover of tasks from Previous E-nights

  1. DONE TCS2019A build - no specific tests needed
    • switch it on during day (after handover), so we can troubleshoot possible problems
    • leave it during night unless new SW problems occurs
  2. ALERT! MODS On-sky checkout - 0.5-1hr - MODS2 deemed offline
    • Quick and dirty test - can we collimate on star before moving to target - never got there due to TCS2019A issues
    • If good seeing, do a MODS dither test with the Pointing Mask to check alignment of detxy y-axis with slit and stage rms (0.75 hr)
      • Stone J good from sunset to 1am
      • Stone K and Stone O until end of night
    • If clear, dither a photometric field from quadrant-to-quadrant.
    • Twilight flats.
  3. ALERT! AZ Science Payback - 2hrs (target sets at 10pm local time - contingent on release of MODS for science)
  4. AGW-1 On-sky checkout - 0.5-1hr
  5. LUCI-1 zeropoint rotator measurement - 0.5 hr
    • Above 30 elevation all night - should do both masks to get enough alignment boxes
      • ID 942616 luci1.8.magmask1.lms 17:56:59 +65 18 42
      • ID 924246 luci1.874.CaT3.lms 14:19:35 +52:49:29
      • Files located at: /home/observer/Service_Observing/E_LUCI_1_ZP_Test
  6. LUCI-1 Rotator Center Check - 0.5 hr - periphery of M92 up all night above elevation limit
    • N30 fieldstop, track mode, 4 PAs (0,180,90,270)
    • Collimate first with an active preset then resend preset in track mode
  7. LUCI-1/AGW-1 transform data collection (take bino data - DX for "free") - 2.5 hours (update PCSconfig with new rotator ZP then take transform - replace old ZP afterwards)
  8. LUCI/LBC mixed-mode - 0.5 hr
  9. LUCI-1/LUCI-2 Non-linearity on-sky tests - 0.5hr - Requires photometric conditions (seeing not critical)
    • RA 17hr (IC 4665)
    • Files located in /home/observer/Service_Observing/E_NonLin_Test (remote room)
  10. D-time Science Programs
    • CO bandhead mergers (3 targets visible across night) - any conditions - filler
    • Located at: /home/observer/Service_Observing/D_Rothberg_Mergers (remote room)
  11. Poor weather backup program - any amount of time - targets cover entire night
    • H-band library - K and M stars
    • H-band library - case example of QSO rest-frame 1.6micron CO-bandhead - filler
    • Located at: /home/observer/Service_Observing/E_Poor_Weather_HbandLibrary (remote room)
  12. Poor weather E task:
    • Examine behavior of open-loop tracking near zenith with MODS.

Nightlog

UT 23:29 (May 22) - ran up both MODS and took sievesnaps - no issues with IMCS lock. Not sure what the issue was now, or what updates were made by RIck in the configuration files. O.Kuhn will investigate offline.

UT 02:30 - P.Kubanek working with J.Williams on update to TCS - note: humidity has been inching upwards and winds are gusting again.

UT 02:49 - tests done - prepping to open - will attempt to point and collimate with MODS near AZ science target - will asess conditions

UT 02:52 - opening enclosure

UT 03:02 - would help if the idiot at the controls put MODS in observe mode and homed the AGWs...

UT 03:12 - found star on left - nothing on right - clouds?

UT 03:23 - testing TCS 2019A - can't seem to find stars...

UT 03:29 - sending collimation preset - failed to find star - reverting to TCS2018B

UT 03:44 - we just hit 95% humidity....closing up

UT 03:51 - going to take the opportuntiy to take MODS cals

Starting with Flats - Dual modsXN.20190523.XXXX

mods1b 0001-0005 u-band

mods1b 0006-0010 g-band

mods1r 0001-0005 r-band

mods1r 0006-0010 i-band

mods1r 0011-0015 z-band

mods2b 0001-0005 u-band

mods2b 0006-0010 g-band

mods2r 0001-0005 r-band

mods2r 0006-0010 i-band

mods2r 0011-0015 z-band

UT 04:35 starting MODS-2 cals for INAF

calmode timed out (calunit) - had to hit retry

UT 04:39 - IMCS lock timed out MODS2 after 120seconds - 2x tries fail on Dual Grating

-testing this on Blue Only Grating - failed, retrying - but script continued on without trying to lock - aborted

UT 04:48 - trying Sievesnap again - worked.. took a long time (we hit 20m/s on winds just as humidity dipped below 90% - OSA checking outside on conditions)

mods2b 0011 - sievesnap

mods2r 0016 - sievesnap

UT 04:53 - trying Dual Grating MODS2 again

-blue went to idle without locking, counts in quad cell remained at bias level - FAILURE

-aborting script - trying to manually lock - blue went to idle quick - Lock-ON button is still orange

red- locked then went to lock requested, then idle

UT 05:07 -instconfig dual grating in command line, then manual imcs lock - blue went to idle with lock-on still in orange

-O.Kuhn running imcsTools - one red cell went to 400 counts, rest bias level

-Failed both sides

-trying old values entered into gui

-Typing Instconfig dual grating in GUI command line - writing down values populated

Blue:

TTFA 26617

TTFB 19269

TTFC 27173

CamFocus 1524

Red:

TTFA 23164

TTFB 17448

TTFC 19286

CamFocus 861

O.Kuhn changing red to 23012,17430,19456 and blue to 26370,19237,27362

-trying manual IMCS lock, no good, bllue went straight to idle

-Trying sievesnap again

-took 1min42 sec for blue 1min49 sec for red, O.Kuhn noted counts really low even when it locked on

mods2b 0012, mods2r 0017 - successfu dua sievesnap

UT 05:30 - mods2b 0013 -Blue Only sievesnap - ran this from script

UT 05:34 - manually configuring instconfig blue imaging then manual click on lock-on

mods2b 0014 - manual sievesnap blue

UT 05:41 - consensus is something is wrong with lasers on MODS-2. Putting both MODS to sleep, chanigng to LUCI

UT 06:10 - configured for Bino-LUCI - winds over 20m/s Humidity 89%

UT 06:21 - init all for both LUCIs

UT 06:22 - taking test darks

luciX.20190523.0001-0002 - 2.51 sec darks

UT 06:44 - humidity 94% - winds still gusting - still closed

UT 06:58 - O.Kuhn noted that todays DATE-OBS header info for MODS files from today and yesterday all have the same values, or groups of the same values, so 10 files would all have the same date and time. We do not see this for LUCI data - taking some LBC biases to see if they have that issue too. All of the MODS data taken on UT 2019-05-23 are in the 20190522 directory in /Repository

-this may be due to the stop/start/stop testing of TCS2019A and TCS2018B - I had to restart IIF connection on LUCI - but the TCS connection may have been lost on MODS - although we did not take any on-sky data

UT 07:06 - 99.9% humidity and the stars have vanished

-after reconnecting tcs - date-obs is corect for MODS data

- also seems reconnect fixed the IMCS lock? but how? instconfig dual grating locked on - but O.Kuhn notes the counts were really low

-trying instconfig blue grating - locked on ~ 1000 counts

-trying red grating - locked on - worked

-taking calibration data for INAF (part deux) mods2N.20190523.XXXX

Biases

mods2b 0015-0019

mods2r 0018-0022

Slitless Pixel Flats Dual

mods2b 0020-0029

mods2r 0023-0027

UT 07:52 - Arc lamps Dual

mods2b 0030-0032

mods2r 0028-0030

UT 08:02 1" slit flats Dual

mods2b 0033-0038

mods2r 0031-0036

UT 08:09 - putting LUCI-2 0.5" mask away - will keep no mask in FPU for now - unless we open.

UT 08:31 - putting MODS to sleep

UT 08:56 - conditions still 99.9% humidity, wind gusts to 20m/s or more, -2.4C ambient temp and cloud around or above us

UT 09:07 - ice on lanai - parking robot MOS LUCI-2. N375 camera blinds mirror cal unit out and no flexure. Ready for exchange.

UT 10:24 - conditions the same - still closed

UT 10:31 - calling end of night - instruments safed


-- BarryRothberg - 22 May 2019
Topic revision: r4 - 23 May 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback